A person brings a mass of 1 kg from infinity to a point A. Initially , the mass was at rest but it moves at a speed of 3 m/s  as it reaches A. The work done by the person on the mass is -5.5J.The gravitational potential at A is ----------------------

Grade:12

1 Answers

vikas askiitian expert
13 years ago

at infinity , V(potential) = 0          

at point A , V = Va

 workdone by gravitational force = mdV = m(Va-0) = mVa

 workdone by person is 5.5J

now , change in kinetic energy = net workdone

     1/2(mv2-0) = mVa - 5.5

   mVa =  4.5+5.5=10

      Va = 10V

thus potential of A is 10V